# Exploiting Style Latent Flows for Generalizing Deepfake Video Detection

### v1

@@ 2 Related work
Deepfake detection. In deepfake detection research, there are two main types: image-level and video-level detectors. The image-level detectors detect fake images by recognizing spatial artifacts of a single frame. A straightforward approach to detect spatial artifacts involves using Xception model [41], a Convolutional Neural Network (CNN) architecture along with MAT [56] that utilizes an attentional mechanism. Face X-ray [28] proposed a method that uses boundaries between forged faces and backgrounds to capture spatial inconsistency. Several methods utilized the frequency-level features to emphasize the artifacts from the image generators [14]. FrePGAN [18] utilized the frequency map transformed from the original RGB image, and LRL [5] integrated RGB images with frequency-aware cues to learn inconsistencies.
In contrast to image-level detectors, video-level detectors take advantage of temporal information by using multiple frames to detect deepfake videos. CNN-GRU [43] captured temporal information from the features extracted by CNNs, and LipForensics [16] utilized the information near the mouth region by employing a network pre-trained on a LipReading dataset. Recently, FTCN [57] directly extracted temporal information using 3D CNNs with a spatial kernel size of $1$. AltFreeze [53] showcases strong generalization capabilities through its unique approach of independently training spatial and temporal information. However, the existing methods have focused on low-level features like pixels, so they ignore the temporal changes of high-level contexts such as facial attributes. To address the problem, we propose a novel approach that considers temporal information in high-level features as well, thereby advancing the state-of-the-art in deepfake detection.
@@ 3.1.4 Style attention module
Method CDF DFD FSh DFo Avg
Xception [7] 73.7 - 72.0 84.5 -
CNN-aug [51] 75.6 - 65.7 74.7 -
PatchForensics [3] 69.6 - 57.8 81.8 -
Multi-task [36] 75.7 - 66.0 77.7 -
FWA [29] 69.5 - 65.5 50.2 -
Two-branch [35] 76.7 - - - -
DCL [45] 82.3 91.6 - - -
Face X-ray [28] 79.5 95.4 92.8 86.8 88.6
NoiseDF [52] 75.9 - - 70.9 -
SLADD [4] 79.7 - - - -
SBI-R50 [44] 85.7 94.0 78.2 91.4 87.3
CNN-GRU [42] 69.8 - 80.8 74.1 -
STIL [15] 75.6 - - - -
LipForensics-s [16] 62.5 - 84.7 84.8 -
RealForensics-s [17] 69.4 - 87.9 89.3 -
FTCN [57] 86.9 94.4 98.8 98.8 94.7
AltFreeze* [53] 89.0 93.7 99.2 99.0 95.2
StyleGRU 62.5 72.2 90.7 77.7 79.8
Ours 89.0 96.1 99.0 99.0 95.8
@@
The output from the SAM is fed into the Temporal Transformer Encoder (TTE) module. The core of TTE is composed of standard transformer encoder blocks [50], with each block including a multi-head self-attention module and multi-layer perceptrons. The results of TTE are the input to the following classification head, yielding the final deepfake prediction through a sigmoid function $\sigma$. Thus,
@@ 4 Experiments
Comparisons. We have compared both image-based and video-based deepfake detection models. Xception [41] is a method of training the Xception model [7], which is a widely used CNN architecture. CNN-aug [51] discovered that CNN-generated images can be readily identified by a CNN model. PatchForensics [3] indicated that the use of patch-based classifiers can contribute to robust deepfake detection. Multi-task [36] approach employed an architecture similar to that of an auto-encoder for deepfake detection. FWA [29] introduced a training data generation technique that relies on deteriorating the quality of the GAN-generated source images. Two-branch [35] method utilized multi-task learning within the context of a deepfake detection dataset. SLADD [4] introduced an algorithm that automatically learns effective augmentation techniques for deepfake detection. DCL [45] applied contrastive learning to the task of deepfake detection in the image domain. NoiseDF [52] compared the features of the cropped face and background squares using noise to detect deepfakes. Face X-ray [28] identified face forgery by segmenting and delineating the blending border between source and target images. CNN-GRU [42] highlighted the potential of utilizing CNN and GRU to investigate temporal features. STIL [15] specifically focused on learning spatio-temporal inconsistency. LipForensics [16] exhibited robust generalization performance through the use of a trained model to capture the representation of natural mouth movement. RealForensics [17] improved the ability to generalize using multimodal architecture. FTCN [57] argued that leveraging temporal inconsistency is beneficial for enhancing the generalizability of deepfake detection algorithms. AltFreeze [53] is an architecture that demonstrates good generalization performance by separately training spatial information and temporal information. For a fair comparison, models trained with datasets other than FF++ are excluded from all experiments, so the performance of both LipForensics and RealForensics is obtained by ‘Scratch’.
4.1 Generalization to cross-datasets
@@
4.2 Generalization to cross-manipulations
For deepfake detection tasks, the evaluation of cross-manipulation performance targets fake videos generated from the same original video but using a different scheme. A common approach is to select three out of the four manipulations provided by FF++, and then the model trained by the three chosen manipulation sets is evaluated by one remaining manipulation set. DF [8] and FS [33] represent identity swap manipulation methods, while F2F [47] and NT [48] denote expression swap manipulation methods. Table 6 presents a comparison of the performance of the cross-manipulation between our model and various models. For expression swap manipulation methods like F2F and NT, they exhibit relatively lower performance compared to Lipforensics [16]. Expression swap inherently lead to inconsistencies in the mouth region, which is why LipForensics appears to perform well. However, our framework displays outstanding performance across all manipulations.
4.3 Ablation Study
4.3.1 Component-wise Ablation Study
@@ 4.4 Analytic Experiments
4.4.2 Qualitative comparison with SAM score
In Fig. 5, a qualitative evaluation is carried out for the SAM scores. Fig. 5(a) illustrates frames extracted from fake clips with notable SAM scores. When considering classification based on low-level temporal cues, it becomes challenging to classify clips demonstrating slow movement. The SAM proposed in our work generates high responses for clips containing slow movement, actively utilizing high-level temporal cues. On the other hand, as shown in Fig. 5(b), for clips with significant motion, it is observed that low responses are generated to focus on low-level temporal cues. This result suggests that the style latent vector flow proposed in this study can be used as a complementary feature to the conventional image-based temporal artifact.
@@ 5 Conclusion
Acknowledgements: This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (2021-0-01341, Artificial Intelligence Graduate School Program(Chung-Ang University) and 2021-0-02067, Next Generation AI for Multi-purpose Video Search) and Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name : Development of high-freedom large-scale user interaction technology using multiple projection spaces to overcome low-light lighting environments. , Project Number : RS-2023-00222280).
@@ 7.1 Style Latent Selection
In Table 5, we can observe that utilizing all latent variables ultimately demonstrates higher generalization performance.
@@
Table 6: Contrastive learning ablation study. We conduct an ablation study on the representation learning methodology. We train on the FF++ dataset and verify performance through video-level AUC scores on the CDF and FSh datasets
7.2 Contrastive learning comparison
Training the StyleGRU in Stage 1 applies a supervised contrastive learning manner using anchor, positive and negative samples with given labels. On the other hand, for representation learning, self-supervised contrasitive learning approaches without providing labels are often employed. We compared the performance difference between our supervised contrastive learning and a self-supervised contrastive learning approach in which clips extracted from the same video as the anchor clip were used as positive clips, and clips from different videos were used as negative clips, without using label.
@@ 8.1 More about style latent variance
8.2 More about Perturbation Robustness

### v2

@@ 2 Related work
Deepfake detection. In deepfake detection research, there are two main types: image-level and video-level detectors. The image-level detectors detect fake images by recognizing spatial artifacts of a single frame. A straightforward approach to detect spatial artifacts involves using Xception model [43], a Convolutional Neural Network (CNN) architecture along with MAT [59] that utilizes an attentional mechanism. Face X-ray [30] proposed a method that uses boundaries between forged faces and backgrounds to capture spatial inconsistency. Several methods utilized the frequency-level features to emphasize the artifacts from the image generators [16]. FrePGAN [20] utilized the frequency map transformed from the original RGB image, and LRL [6] integrated RGB images with frequency-aware cues to learn inconsistencies. Recently, algorithms utilize methods including blending artifacts [46, 5, 14, 2] or separating elements relevant to detection from irrelevant ones [14, 58] during their training. Image-level detectors demonstrate high generalization performance. However, image-based methods have the limitation of not utilizing the temporal cues that inevitably occur during the generation of deepfake videos.
In contrast to image-level detectors, video-level detectors take advantage of temporal information by using multiple frames to detect deepfake videos. CNN-GRU [45] captured temporal information from the features extracted by CNNs, and LipForensics [18] utilized the information near the mouth region by employing a network pre-trained on a LipReading dataset. Recently, FTCN [60] directly extracted temporal information using 3D CNNs with a spatial kernel size of $1$. AltFreeze [55] showcases strong generalization capabilities through its unique approach of independently training spatial and temporal information. However, the existing methods have focused on low-level features like pixels, so they ignore the temporal changes of high-level contexts such as facial attributes. To improve detection capabilities, we propose a novel approach that considers temporal features at both low-level and high-level, thereby enhancing the overall context and advancing the state-of-the-art in deepfake detection.
@@ 3.1.4 Style attention module
where $s(\bullet)$ represents a softmax function.
@@
The output from the SAM is linearly projected into 16 time steps (n=16) and then fed into the Temporal Transformer Encoder (TTE) module. The core of TTE is composed of standard transformer encoder blocks [52], with each block including a multi-head self-attention module and multi-layer perceptrons. The results of TTE are the input to the following classification head, yielding the final deepfake prediction through a sigmoid function $\sigma$. Thus,
@@ 4 Experiments
Method CDF DFD FSh DFo Avg
Xception [8] 73.7 - 72.0 84.5 -
CNN-aug [53] 75.6 - 65.7 74.7 -
PatchForensics [4] 69.6 - 57.8 81.8 -
Multi-task [38] 75.7 - 66.0 77.7 -
FWA [31] 69.5 - 65.5 50.2 -
Two-branch [37] 76.7 - - - -
DCL [47] 82.3 91.6 - - -
Face X-ray [30] 79.5 95.4 92.8 86.8 88.6
NoiseDF [54] 75.9 - - 70.9 -
SLADD [5] 79.7 - - - -
SBI-R50 [46] 85.7 94.0 78.2 91.4 87.3
CNN-GRU [44] 69.8 - 80.8 74.1 -
STIL [17] 75.6 - - - -
LipForensics [18] 82.4 - 97.1 97.6 -
RealForensics* [19] 85.6 82.2 99.6 99.8 91.8
FTCN* [60] 86.9 94.4 98.8 98.8 94.7
AltFreeze* [55] 89.0 93.7 99.2 99.0 95.2
StyleGRU 62.5 72.2 90.7 77.7 79.8
Ours 89.0 96.1 99.0 99.0 95.8
Comparisons. We have compared both image-based and video-based deepfake detection models. Xception [43] is a method of training the Xception model [8], which is a widely used CNN architecture. CNN-aug [53] discovered that CNN-generated images can be readily identified by a CNN model. PatchForensics [4] indicated that the use of patch-based classifiers can contribute to robust deepfake detection. Multi-task [38] approach employed an architecture similar to that of an auto-encoder for deepfake detection. FWA [31] introduced a training data generation technique that relies on deteriorating the quality of the GAN-generated source images. Two-branch [37] method utilized multi-task learning within the context of a deepfake detection dataset. SLADD [5] introduced an algorithm that automatically learns effective augmentation techniques for deepfake detection. SBI [46] proposed a novel approach to blending artifacts using self-images. DCL [47] applied contrastive learning to the task of deepfake detection in the image domain. NoiseDF [54] compared the features of the cropped face and background squares using noise to detect deepfakes. Face X-ray [30] identified face forgery by segmenting and delineating the blending border between source and target images. CNN-GRU [44] highlighted the potential of utilizing CNN and GRU to investigate temporal features. STIL [17] specifically focused on learning spatio-temporal inconsistency. LipForensics [18] exhibited robust generalization performance through the use of a trained model to capture the representation of natural mouth movement. RealForensics [19] improved the ability to generalize using multimodal architecture. FTCN [60] argued that leveraging temporal inconsistency is beneficial for enhancing the generalizability of deepfake detection algorithms. AltFreeze [55] is an architecture that demonstrates good generalization performance by separately training spatial information and temporal information.
4.1 Generalization to cross-datasets
@@
4.2 Generalization to cross-manipulations
For deepfake detection tasks, the evaluation of cross-manipulation performance targets fake videos generated from the same original video but using a different scheme. A common approach is to select three out of the four manipulations provided by FF++, and then the model trained by the three chosen manipulation sets is evaluated by one remaining manipulation set. DF [9] and FS [35] represent identity swap manipulation methods, while F2F [49] and NT [50] denote expression swap manipulation methods. AltFreeze [55] does not officially provide weights for this experiment setting. Therefore, it is not possible for us to reproduce. Table 2 presents a comparison of the performance of the cross-manipulation between our model and various models. For expression swap manipulation methods like F2F and NT, they exhibit relatively lower performance compared to Lipforensics [18]. Expression swap inherently lead to inconsistencies in the mouth region, which is why LipForensics appears to perform well. However, our framework displays outstanding performance across all manipulations.
4.3 Ablation Study
4.3.1 Component-wise Ablation Study
@@ 5 Conclusion
Acknowledgements: This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (2021-0-01341, Artificial Intelligence Graduate School Program(Chung-Ang University), 2020-0-01336, AIGS program (UNIST), and 2021-0-01778 Development of human image synthesis and discrimination technology below the perceptual threshold) and Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name : Development of high-freedom large-scale user interaction technology using multiple projection spaces to overcome low-light lighting environments. , Project Number : RS-2023-00222280).
[…]
@@ 7.1 Style Latent Selection
Metric CDF Fsh
No differencing 88.4 98.4
2nd-order differencing 88.8 98.8
Ours(1st-order differencing) 89.0 99.0
Table 6: Flow modeling metric comparison experiment.
As presented in Table 6, it reveals that the StyleGRU achieves superior performance when utilizing a metric based on first-order differencing, compared to other approaches such as neglecting differencing or employing second-order differencing. Nonetheless, the small performance gaps suggest that the GRU layer lets the style feature independent of the metric.
@@
Table 7: Contrastive learning ablation study.
@@
7.2 Contrastive learning comparison
Training the StyleGRU in Stage 1 applies a supervised contrastive learning manner using anchor, positive and negative samples with given labels. On the other hand, for representation learning, self-supervised contrasitive learning approaches without providing labels are often employed. We compared the performance difference between our supervised contrastive learning and a self-supervised contrastive learning approach in which clips extracted from the same video as the anchor clip were used as positive clips, and clips from different videos were used as negative clips, without using label. We conduct an ablation study on the representation learning methodology. We train on the FF++ dataset and verify performance through video-level AUC scores on the CDF and FSh datasets.
7.3 More about Perturbation Robustness
Method raw c23 c40
FTCN* 99.5 99.0 70.2
Ours 99.6 99.1 71.8
Table 9: Experiment on low-quality images.
As shown in Table 9, we performed an experiment to verify if our algorithm generalizes well to low-quality samples even when trained on high-quality ones. In the results, it can be observed that our method exhibits relatively robust performance against low-quality samples.
Method CDF DFD KoDF Avg
FTCN* 86.9 94.4 69.4 83.6
Altfreeze* 89.0 93.7 69.8 84.2
Ours 89.0 96.1 71.1 85.4
Table 10: Generalization to distinct datasets from training data.
@@ 8.1 More about style latent variance
8.2 Qualitative comparison with SAM score
In Figure 7, a qualitative evaluation is carried out for the SAM scores. Figure 7(a) illustrates frames extracted from fake clips with notable SAM scores. When considering classification based on low-level temporal cues, it becomes challenging to classify clips demonstrating slow movement. The SAM proposed in our work generates high responses for clips containing slow movement, actively utilizing high-level temporal cues. On the other hand, as shown in Figure 7(b), for clips with significant motion, it is observed that low responses are generated to focus on low-level temporal cues. This result suggests that the style latent vector flow proposed in this study can be used as a complementary feature to the conventional image-based temporal artifact.